OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

look at anthony's question. i answered one just like this one.

OpenStudy (lgbasallote):

hint: use this formula \[\large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\] are you familiar with this formula?

hero (hero):

Input each point separately into the formula y = mx + b

OpenStudy (anonymous):

@hero but how

OpenStudy (anonymous):

@lgbasallote nope